OpenStudy (andrewkaiser333):

Diane delivers newspapers for $5 a day plus $0.04 per newspaper delivered. Jeremy delivers for $2 a day plus $0.10 per newspaper delivered how many are needed to get them to have the same amount of money earned.

OpenStudy (andrewkaiser333):

@jdoe0001

OpenStudy (andrewkaiser333):

@keke_luvs_u

OpenStudy (austinl):

Okay, you need to set up two different equations. \(\Large{\text{Diane}=5+0.04x}\) \(\Large{\text{Jeremy}=2+0.10x}\) Where x is number of papers delivered. You then need to set these equal and solve for x.

OpenStudy (andrewkaiser333):

k

OpenStudy (andrewkaiser333):

50=x ?????

OpenStudy (austinl):

Indeed! Good job!

OpenStudy (andrewkaiser333):

thanks

OpenStudy (andrewkaiser333):

got one more

OpenStudy (andrewkaiser333):

a student store sold a total of 55 shirts for $620. the shirts sold were either red or white. red sold for $12 each. white sold for $10 each. how many of each color were sold???

OpenStudy (austinl):

x=red y=white \(x+y=55\) \(12x+10y=620\) Solve for the variables.

OpenStudy (andrewkaiser333):

62=y 51.666666666= x

OpenStudy (andrewkaiser333):

????

OpenStudy (austinl):

No decimal answers, leave in exact form.

OpenStudy (andrewkaiser333):

so 52

OpenStudy (andrewkaiser333):

?????

OpenStudy (austinl):

Not the answer that I got at all actually.

OpenStudy (andrewkaiser333):

how you do it?

OpenStudy (austinl):

\(x+y=55\) \(x=55-y\) \(12x+10y=620\) \(12(55-y)+10y=620\) \(660-2y=620\) \(-2y=-40\) \(\color{red}{y=20}\) \(x+20=55\) \(\color{red}{x=35}\)

OpenStudy (andrewkaiser333):

k

OpenStudy (andrewkaiser333):

i see thanks

OpenStudy (austinl):

No problem.